# Investigating Adversarial Attacks in Software Analytics via Machine Learning Explainability

### v1

@@ 4.2. Present Experimental Results
Comparison with the baselines: We compare our approach with four state-of-the-art techniques: Zoo (Chen et al., 2017b), Boundary attack (Brendel et al., 2017), PermuteAttack (Hashemi and Fathi, 2020), and HopSkipJump (Chen et al., 2020), in terms of ASR metric values and imperceptibility (e.g., minimal $\ell_{0}$ perturbation) across various ML models. It is important to note that we consider the best result (highlighted in bold text) from the three explanation-guided attacks for comparison with the baselines. Table 5 demonstrates that our technique outperforms HopSkipJump and Boundary attacks significantly in terms of imperceptibility. In many cases (highlighted in bold text), we observe that our approach outperforms HopSkipJump and Boundary attacks both in ASR metric values and imperceptibility. Additionally, HopSkipJump and Boundary attacks modify all features of test instances, thereby contradicting the imperceptibility property of the adversarial examples defined in Section 3.4. Shifting to the Zoo attack, in the majority of cases (highlighted in bold text), our approach surpasses the Zoo attack in terms of ASR metric values and imperceptibility. Furthermore, the Zoo attack fails to generate a single adversarial example for a few models (highlighted in grey cells). Moreover, regarding imperceptibility, our approach exhibits superior performance over the Zoo attack.
@@
Table 7 presents the ASR metric and top-$k$ values for both the explanation-guided and selected state-of-the-art adversarial attacks using the balanced test data. From Table 7, it is evident that our explanation-guided adversarial attack approach outperforms the ZOO, HopSkipJump, and Boundary attack techniques in terms of the ASR metric and imperceptibility. Although the PermuteAttack performs well in terms of the ASR metric, our approach achieves a better balance between the ASR metric and imperceptibility. Additionally, the PermuteAttack fails to generate a single adversarial example for the ADA model trained on the Java project dataset and performs poorly on the same model trained on the Postgres dataset. Therefore, we reach a similar conclusion even after balancing the test dataset.
@@
Figure 6. Comparison of our approach with the makeshift tools regarding ASR metric. The first row compares our approach with the makeshift tools when we change the top-k important features identified by SHAP (SH, Fig. 6(a)), LIME (LM. Fig. 6(b)), and PyExplainer (PE, Fig. 6(c)) for the CLCDSA dataset. Similarly, the second, third, and fourth rows compare the code review, Postgres, and BCB datasets, respectively. Note: ASR metric values are rounded for better visualization.
@@
Figure 7. Comparison of our approach with the makeshift tools regarding ASR metric for the balanced test data. The first and second rows compare our approach with the makeshift tools when we change the top-k important features identified by SHAP, LIME, and PyExplainer for the Java project and Postgres datasets. Note: ASR metric values are rounded for better visualization.
@@ 5. Threats to Validity
5.1. Internal Threat
The first internal threat is the accuracy of ML models. However, we countered this threat by choosing the hyperparameter settings described in previous studies (Catolino et al., 2019; Pornprasit et al., 2021). We used SMOTE and Autospearman and applied grid search, random search, and Bayesian optimization to find the best hyperparameter combinations for each ML model. Thus, our selected ML models have achieved good accuracy. We addressed the potential threat concerning the quality of important features identified by ML explainability techniques by choosing two widely recognized model-agnostic methods, SHAP and LIME. Additionally, we used PyExplainer (Pornprasit et al., 2021), explicitly designed to explain JIT defect models. Therefore, the important features found by these three techniques are good enough to generate adversarial examples. Another internal threat could be selecting the top-$k$ features to generate adversarial examples. However, Section 3.4 describes how we mitigated this threat. Our experiments show that LIME and PyExplainer sometimes generate different feature importance rankings across multiple executions of the same instance. This variability in experimental outcomes is not a limitation of our study.
5.2. Construct Threat
One potential threat could be how changes to top-$k$ features might affect source code. However, we addressed this concern by considering the characteristics of black-box adversarial attacks performed on extracted features stored in tabular format. Similar to studies in the financial domain (Cartella et al., 2021; Mathov et al., 2020; Du et al., 2023) and in the software analytics (Zhang et al., 2021), we assume access to both training and test datasets for black-box attacks. We constructed surrogate ML models to carry out adversarial attacks, following the approaches outlined in (Cartella et al., 2021; Mathov et al., 2020). In addition, similar to studies in computer vision (Xu et al., 2020; Inkawhich et al., 2019; Simonetto et al., 2021), we attack the feature-space (e.g., metrics or properties of source code in tabular format) while generating adversarial examples. Since we only have access to the extracted features in tabular format (e.g., training and testing data) rather than the source code itself, and considering that these extracted features serve as the direct input to ML models during training, we explicitly state that observing the impact of altered features on the source code falls outside the scope of this study.
@@ 5.2. Construct Threat
5.3. External Threat
The generalizability of our experimental results might threaten the external validity. However, we resolved this threat by choosing seven classical ML models extensively studied in different software analytics tasks, three ML explainability techniques and six distinct datasets used in the previous studies (Feng et al., 2024; Wu et al., 2022; Hu et al., 2022; Catolino et al., 2019; Roy et al., 2022; Pornprasit et al., 2021; Chen et al., 2017a). Thus, we considered $6\times 3\times 7=126$ experimental combinations for the analysis. Therefore, our extensive experimental findings are possibly sufficiently robust to generalize our study.
6. Related Work
@@
Szegedy et al. (Szegedy et al., 2013) first introduced the concept of adversarial attack for image classifiers. To the best of our knowledge, we have not found any research on adversarial attacks targeting classical ML models trained on extracted features stored in tabular format within software analytics tasks, such as JIT defect prediction, clone detection, and classification of useful code review comments. Several studies have been done on adversarial attacks targeting ML models trained on tabular data in other domains (Ballet et al., 2019; Hashemi and Fathi, 2020; Grosse et al., 2017; Papernot et al., 2016; Mathov et al., 2020; Esmaeilpour et al., 2022; Kireev et al., 2022; Melo et al., 2023; Hyeong et al., 2022; Alshantti et al., 2023; Simonetto et al., 2023; Pleiter et al., 2023; Hu et al., 2021). Ballet et al. (Ballet et al., 2019) were the first to introduce systematic adversarial attacks on fully connected neural networks trained on tabular financial data. They altered feature values and modified less significant features based on human judgment to craft imperceptible adversarial examples. Similar approaches have been proposed by Grosse et al. (Grosse et al., 2017), Hashemi et al. (Hashemi and Fathi, 2020), Levy et al. (Mathov et al., 2020), and Papernot et al. (Papernot et al., 2016). Additionally, Cartella et al. (Cartella et al., 2021) adapted state-of-the-art adversarial attack techniques such as the Zoo attack (Chen et al., 2017b), boundary attack (Brendel et al., 2017), and HopSkipJump attack (Chen et al., 2020) from computer vision to the financial domain using tabular data. A different study by Levy et al. (Mathov et al., 2020) introduced an approach for transforming a surrogate model while preserving the characteristics of the original model, enabling the application of established generation techniques. They demonstrated that a slight human imperceptible perturbation of the input image could change the prediction of the image classifiers. All of the aforementioned studies considered various constraints when generating adversarial examples, none of which are applicable to tabular data in software analytics tasks.
@@
A plethora of research has been conducted on adversarial attacks targeting various software analytics tasks. Liu et al. (Liu et al., 2021), and Erwin et al. (Quiring et al., 2019) proposed a practical black-box attack on source code authorship identification classifiers based on a set of semantically equivalent program transformations. Nguyen et al. (Nguyen et al., 2021) showed that state-of-the-art API recommender systems are vulnerable to adversarial attacks if attackers corrupt the training corpus by injecting malicious data. Chen et al. (Chen et al., 2017a) developed the EvenAttack model considering the contributions of the features of instances to the malware detection problem. Liu et al. (Liu et al., 2019) proposed a novel method (ATMPA) considering the gradient descent and L-norm optimization method. An attacker can use their technique to generate adversarial examples by introducing tiny perturbations to the input data. Grosse et al. and Suciu et al. (Suciu et al., 2019) demonstrated how adversarial examples fool the malware detection classifiers with slightly modified input data. However, none of the methods proposed for adversarial attacks on ML models in software analytics consider the importance of the features identified by ML explainability techniques. Our work aims to manipulate the important features identified by ML explainability techniques to generate adversarial examples. Moreover, adversarial attacks in the feature-space have been studied in computer vision (Xu et al., 2020; Inkawhich et al., 2019; Simonetto et al., 2021). However, this area remains largely unexplored in software analytics tasks. Therefore, our work aims to advance research in this direction.
@@
Severi et al. (Severi et al., 2021) applied ML explainability to select important features and values for malware classifier adversarial attacks. Unlike their work, our approach is designed specifically for ML inference, not training. Furthermore, our methodology diverges in selecting and modifying the top-$k$ important features while generating adversarial examples. Amich et al. (Amich and Eshete, 2021, 2022) employ ML explainability to enhance and diagnose evasion attacks on ML models. Their study focuses on ML models trained on image classifier datasets like MNIST and CIFAR-10, with a different approach to perturbation and adversarial example generation than ours. Zhang et al. (Zhang et al., 2024) proposed a novel ensemble-based adversarial attack approach that focuses on balancing two key aspects: 1) transferability and 2) imperceptibility, based on model interpretability for image data. Sun et al. (Sun et al., 2023b) proposed an explainability-guided, model-agnostic testing framework to assess the robustness of malware detectors through feature-space manipulation. The framework uses Accured Malicious Magnitude (AMM) to identify the fragile features for manipulation while generating adversarial examples. The primary distinction between our work and that of Sun et al. (Sun et al., 2023b) is that we introduced the concept of the Reverse Elbow Method for identifying fragile features (e.g., top-$k$ important features) for feature-space manipulation. Another important distinction is that their adversarial attack technique is a white-box method, assuming the attacker has complete knowledge of the model architecture and training dataset. In contrast, our explanation-guided adversarial attack technique does not have access to any information about the model’s architecture. Additionally, we applied a multi-objective optimization function, as defined in Section 3.4, to generate adversarial examples.
@@
Numerous studies have investigated adversarial attacks targeting deep learning models and large language models (LLMs) trained on source code for various software analytics tasks such as clone detection, source code authorship attribution, vulnerability detection, code comment generation, and code summarization (Yefet et al., 2020; Srikant et al., 2021; Zhang et al., 2020; Tian et al., 2021; Zhou et al., 2022; Bielik and Vechev, 2020; Springer et al., 2020; Li et al., 2022, 2021; Tian et al., 2023; Zhou et al., 2024; Sun et al., 2023a; Wan et al., 2022; Zhang et al., 2021; Wong et al., 2022; Liu and Zhang, 2024; Jia et al., 2023; Zhu and Zhang, 2023; He and Vechev, 2023). Yefet et al. (Yefet et al., 2020) proposed a white-box attack technique, DAMP, that changes the identifier’s name in the code snippet based on the gradient information of the victim model. Srikant et al. (Srikant et al., 2021) generated adversarial examples for models of code using optimized obfuscation. Zhang et al. (Zhang et al., 2020) proposed a black-box attack technique MHM based on Metropolis-Hasting sampling-based variable renaming approach. Tian et al. (Tian et al., 2021) introduced QMDP (Q-Learning-based Markov decision process) to generate adversarial examples for models of code based on the semantic equivalent program transformations. Their experimental results demonstrated that QMDP generates adversarial examples and enhances the robustness of the source code classification models by over 44%.
@@
Pour et al. (Pour et al., 2021) developed a search-based adversarial test generation framework to measure the robustness of the neural source code embedding methods (i.e., Code2vec, Code2seq, and CodeBERT). Their experimental results show that the generated adversarial examples can, on average, decrease the performance of these embedding methods from 5.41% to 9.58%. Applis et al. (Applis et al., 2021) proposed a testing framework: LAMPION, to assess the robustness of the ML-based program analysis tools adopting metamorphic program transformations. Zhou et al. (Zhou et al., 2022) studied the robustness of the code comment generation tasks in adversarial settings. They proposed ACCENT to generate adversarial examples by substituting identifiers in code snippets to generate syntactically correct and semantically similar ones to original code snippets. Recently, several works have focused on assessing the robustness of pre-trained code models under adversarial attacks (Yang et al., 2022; Du et al., 2023; Zeng et al., 2022; Jha and Reddy, 2023). Yang et al. (Yang et al., 2022) performed a natural attack on the pre-trained code models. They transformed the code snippet, preserving the original inputs’ operational and natural semantics. Bielik et al. (Bielik and Vechev, 2020), and Springer et al. (Springer et al., 2020) both evaluated the robustness of models of code in adversarial settings. Schuster et al. (Schuster et al., 2021) demonstrated that carefully designed adversarial files attached to the training corpus make the code completion model vulnerable.

### v2

@@ 4.2 Present Experimental Results
Comparison with the baselines: We compare our approach with four state-of-the-art techniques: Zoo [59], Boundary attack [60], PermuteAttack [61], and HopSkipJump [62], in terms of ASR metric values and imperceptibility (e.g., minimal $\ell_{0}$ perturbation) across various ML models. It is important to note that we consider the best result (highlighted in bold text) from the three explanation-guided attacks for comparison with the baselines. Table 7 demonstrates that our technique outperforms HopSkipJump and Boundary attacks significantly in terms of imperceptibility. In many cases (highlighted in bold text), we observe that our approach outperforms HopSkipJump and Boundary attacks both in ASR metric values and imperceptibility. Additionally, HopSkipJump and Boundary attacks modify all features of test instances, thereby contradicting the imperceptibility property of the adversarial examples defined in Section 3.4. Shifting to the Zoo attack, in the majority of cases (highlighted in bold text), our approach surpasses the Zoo attack in terms of ASR metric values and imperceptibility. Furthermore, the Zoo attack fails to generate a single adversarial example for a few models (highlighted in grey cells). Moreover, regarding imperceptibility, our approach exhibits superior performance over the Zoo attack.
Table 9: Attack Success Rate (ASR) metric values and the Top-$k$ modified features (e.g., a minimal $\ell_{0}$ perturbation) when we apply explanation-guided adversarial attacks and selected baseline attacks on the ML models for the balanced Java project and Postgres datasets
Java Project SHAP LIME PyExplainer PermuteAttack Zoo Boundary HopSkipJump ML Model ASR Top-$k$ ASR Top-$k$ ASR Top-$k$ ASR Top-$k$ ASR Top-$k$ ASR Top-$k$ LR 40.98 Top-3 48.78 Top-3 25.06 97.98 Top-7 23.38 Top-13 45.43 Top-27 44.98 Top-27 DT 60.71 Top-2 47.49 Top-1 20.59 97.95 Top-5 0.0 Top-0 55.12 Top-27 55.12 Top-27 RF 44.58 Top-2 39.58 Top-3 9.07 98.75 Top-6 0.0 Top-0 100 Top-27 100 Top-27 MLP 50.09 Top-2 38.4 Top-1 22.26 99.27 Top-7 16.67 Top-12 45.82 Top-27 45.08 Top-27 ADA 46.62 Top-3 41.59 Top-3 19.11 0.0 Top-0 0.0 Top-0 46.61 Top-27 46.42 Top-27 BAG 51.96 Top-2 44.21 Top-2 17.86 56.19 Top-6 0.02 Top-1 99.63 Top-27 99.63 Top-27 GBC 55.97 Top-3 46.27 Top-1 14.37 99.25 Top-6 0.0 Top-0 41.6 Top-27 41.6 Top-27 Postgres LR 85.89 Top-1 83.06 Top-3 34.54 99.46 Top-6 99.32 Top-8 41.8 Top-12 41.8 Top-12 DT 85.63 Top-2 82.92 Top-2 6.9 85.87 Top-4 85.87 Top-12 66.54 Top-12 68.55 Top-12 RF 65.45 Top-2 72.91 Top-2 21.36 100 Top-4 91.99 Top-4 86.08 Top-12 84.77 Top-12 MLP 56.99 Top-1 59.45 Top-2 22.4 87.94 Top-6 88.08 Top-11 48.9 Top-12 48.76 Top-12 ADA 84.09 Top-3 56.98 Top-1 19.61 8.34 Top-3 56.97 Top-4 40.93 Top-12 40.93 Top-12 BAG 83.59 Top-2 79.56 Top-3 36.98 100 Top-4 86.71 Top-7 50.13 Top-12 49.47 Top-12 GBC 70.99 Top-2 73.54 Top-2 19.75 100 Top-4 85.69 Top-4 93.24 Top-12 92.05 Top-12
Note: Bolded cells indicate cases where the attack techniques failed to produce any successful adversarial examples
@@
Table 9 presents the ASR metric and top-$k$ values for both the explanation-guided and selected state-of-the-art adversarial attacks using the balanced test data. The bolded cells highlighted in grey represent cases where the attack techniques failed to produce any successful adversarial examples. From Table 9, it is evident that our explanation-guided adversarial attack approach outperforms the ZOO, HopSkipJump, and Boundary attack techniques in terms of the ASR metric and imperceptibility. Although the PermuteAttack performs well in terms of the ASR metric, our approach achieves a better balance between the ASR metric and imperceptibility. Additionally, the PermuteAttack fails to generate a single adversarial example for the ADA model trained on the Java project dataset and performs poorly on the same model trained on the Postgres dataset. Therefore, we reach a similar conclusion even after balancing the test dataset.
While the tabular results provide detailed comparisons, we also employ win-rate analysis to more effectively summarize the overall performance of our attack approach relative to baseline methods. Win-rate analysis offers a straightforward yet powerful means of comparing two approaches across multiple experimental settings (e.g., different datasets or models). We evaluate whether the proposed method outperforms, matches, or underperforms for each setting relative to the baseline. The win-rate is computed as the proportion of settings where the proposed method outperforms the baseline, defined as:
$\text{Win Rate}_{A\text{ vs }B}=\frac{\text{Number of Wins of }A}{\text{Total Number of Comparisons}}\times 100\%$ (5)
Where, a win is counted if method $A$ (e.g., our proposed approach) achieves better performance than method $B$ (e.g., baselines) according to the evaluation metric of interest (e.g., ASR).
Figure 6 compares the win-rates between the explanation-guided attack and four baseline adversarial attacks— PermuteAttack, Zoo, Boundary, and HopSkipJump. In each pairwise comparison, the explanation-guided attack consistently achieves a higher win-rate than the corresponding baseline methods regarding the imperceptibility metric. However, although the PermuteAttack performs better than our approach, considering the ASR metric, our approach outperforms the other three baseline attacks. Overall, the win-rate analysis suggests that the explanation-guided attack is generally more effective in generating successful adversarial examples than the classical baseline attacks. The consistent improvement across multiple baselines highlights the robustness and reliability of the proposed attack strategy.
(a) ASR on Unbalanced Data
(b) ASR on Balanced Data
(c) Imperceptibility on Unbalanced Data
(d) Imperceptibility on Balanced Data
Figure 6: Comparison of the explanation-guided attack with four baselines based on ASR and imperceptibility metrics. Each group illustrates the performance difference between the proposed method and each baseline method.
@@
Figure 7: Comparison of our approach with the makeshift tools regarding ASR metric. The first row compares our approach with the makeshift tools when we change the top-k important features identified by SHAP (SH, Fig. 7(a)), LIME (LM. Fig. 7(b)), and PyExplainer (PE, Fig. 7(c)) for the CLCDSA dataset. Similarly, the second, third, and fourth rows compare the code review, Postgres, and BCB datasets, respectively. Note: ASR metric values are rounded for better visualization.
@@
Figure 8: Comparison of our approach with the makeshift tools regarding ASR metric for the balanced test data. The first and second rows compare our approach with the makeshift tools when we change the top-k important features identified by SHAP, LIME, and PyExplainer for the Java project and Postgres datasets. Note: ASR metric values are rounded for better visualization.
@@ 5 Threats to Validity
5.1 Internal Threat
The first internal threat is the accuracy of ML models. However, we countered this threat by choosing the hyperparameter settings described in previous studies [13, 35]. We used SMOTE and Autospearman and applied grid search, random search, and Bayesian optimization to find the best hyperparameter combinations for each ML model. Thus, our selected ML models have achieved good accuracy. We addressed the potential threat concerning the quality of important features identified by ML explainability techniques by choosing two widely recognized model-agnostic methods, SHAP and LIME. Additionally, we used PyExplainer [35], explicitly designed to explain JIT defect models. Therefore, the important features found by these three techniques are good enough to generate adversarial examples. Another internal threat could be selecting the top-$k$ features to generate adversarial examples. However, Section 3.4 describes how we mitigated this threat. Our experiments show that LIME and PyExplainer sometimes generate different feature importance rankings across multiple executions of the same instance. This variability in experimental outcomes is not a limitation of our study.
5.2 Construct Threat
One potential threat could be how changes to top-$k$ features might affect source code. However, we addressed this concern by considering the characteristics of black-box adversarial attacks performed on extracted features stored in tabular format. Similar to studies in the financial domain [81, 76, 31] and in the software analytics [85], we assume access to both training and test datasets for black-box attacks. We constructed surrogate ML models to carry out adversarial attacks, following the approaches outlined in [81, 76]. In addition, similar to studies in computer vision [52, 53, 54], we attack the feature-space (e.g., metrics or properties of source code in tabular format) while generating adversarial examples. Since we only have access to the extracted features in tabular format (e.g., training and testing data) rather than the source code itself, and considering that these extracted features serve as the direct input to ML models during training, we explicitly state that observing the impact of altered features on the source code falls outside the scope of this study.
@@ 5.2 Construct Threat
5.3 External Threat
The generalizability of our experimental results might threaten the external validity. However, we resolved this threat by choosing seven classical ML models extensively studied in different software analytics tasks, three ML explainability techniques and six distinct datasets used in the previous studies [2, 4, 3, 13, 65, 35, 21]. Thus, we considered $6\times 3\times 7=126$ experimental combinations for the analysis. Therefore, our extensive experimental findings are possibly sufficiently robust to generalize our study.
6 Related Work
@@
Szegedy et al. [18] first introduced the concept of adversarial attack for image classifiers. To the best of our knowledge, we have not found any research on adversarial attacks targeting classical ML models trained on extracted features stored in tabular format within software analytics tasks, such as JIT defect prediction, clone detection, and classification of useful code review comments. Several studies have been done on adversarial attacks targeting ML models trained on tabular data in other domains [80, 61, 88, 89, 76, 90, 91, 92, 93, 94, 95, 96, 97]. Ballet et al. [80] were the first to introduce systematic adversarial attacks on fully connected neural networks trained on tabular financial data. They altered feature values and modified less significant features based on human judgment to craft imperceptible adversarial examples. Similar approaches have been proposed by Grosse et al. [88], Hashemi et al. [61], Levy et al. [76], and Papernot et al. [89]. Additionally, Cartella et al. [81] adapted state-of-the-art adversarial attack techniques such as the Zoo attack [59], boundary attack [60], and HopSkipJump attack [62] from computer vision to the financial domain using tabular data. A different study by Levy et al. [76] introduced an approach for transforming a surrogate model while preserving the characteristics of the original model, enabling the application of established generation techniques. They demonstrated that a slight human imperceptible perturbation of the input image could change the prediction of the image classifiers. All of the aforementioned studies considered various constraints when generating adversarial examples, none of which are applicable to tabular data in software analytics tasks.
@@
A plethora of research has been conducted on adversarial attacks targeting various software analytics tasks. Liu et al. [19], and Erwin et al. [20] proposed a practical black-box attack on source code authorship identification classifiers based on a set of semantically equivalent program transformations. Nguyen et al. [29] showed that state-of-the-art API recommender systems are vulnerable to adversarial attacks if attackers corrupt the training corpus by injecting malicious data. Chen et al. [21] developed the EvenAttack model considering the contributions of the features of instances to the malware detection problem. Liu et al. [34] proposed a novel method (ATMPA) considering the gradient descent and L-norm optimization method. An attacker can use their technique to generate adversarial examples by introducing tiny perturbations to the input data. Grosse et al. and Suciu et al. [33] demonstrated how adversarial examples fool the malware detection classifiers with slightly modified input data. However, none of the methods proposed for adversarial attacks on ML models in software analytics consider the importance of the features identified by ML explainability techniques. Our work aims to manipulate the important features identified by ML explainability techniques to generate adversarial examples. Moreover, adversarial attacks in the feature-space have been studied in computer vision [52, 53, 54]. However, this area remains largely unexplored in software analytics tasks. Therefore, our work aims to advance research in this direction.
@@
Severi et al. [28] applied ML explainability to select important features and values for malware classifier adversarial attacks. Unlike their work, our approach is designed specifically for ML inference, not training. Furthermore, our methodology diverges in selecting and modifying the top-$k$ important features while generating adversarial examples. Amich et al. [55, 56] employ ML explainability to enhance and diagnose evasion attacks on ML models. Their study focuses on ML models trained on image classifier datasets like MNIST and CIFAR-10, with a different approach to perturbation and adversarial example generation than ours. Zhang et al. [98] proposed a novel ensemble-based adversarial attack approach that focuses on balancing two key aspects: 1) transferability and 2) imperceptibility, based on model interpretability for image data. Sun et al. [99] proposed an explainability-guided, model-agnostic testing framework to assess the robustness of malware detectors through feature-space manipulation. The framework uses Accured Malicious Magnitude (AMM) to identify the fragile features for manipulation while generating adversarial examples. The primary distinction between our work and that of Sun et al. [99] is that we introduced the concept of the Reverse Elbow Method for identifying fragile features (e.g., top-$k$ important features) for feature-space manipulation. Another important distinction is that their adversarial attack technique is a white-box method, assuming the attacker has complete knowledge of the model architecture and training dataset. In contrast, our explanation-guided adversarial attack technique does not have access to any information about the model’s architecture. Additionally, we applied a multi-objective optimization function, as defined in Section 3.4, to generate adversarial examples.
@@
Numerous studies have investigated adversarial attacks targeting deep learning models and large language models (LLMs) trained on source code for various software analytics tasks such as clone detection, source code authorship attribution, vulnerability detection, code comment generation, and code summarization [100, 101, 27, 102, 26, 22, 23, 103, 104, 105, 106, 107, 108, 85, 109, 110, 111, 112, 113]. Yefet et al. [100] proposed a white-box attack technique, DAMP, that changes the identifier’s name in the code snippet based on the gradient information of the victim model. Srikant et al. [101] generated adversarial examples for models of code using optimized obfuscation. Zhang et al. [27] proposed a black-box attack technique MHM based on Metropolis-Hasting sampling-based variable renaming approach. Tian et al. [102] introduced QMDP (Q-Learning-based Markov decision process) to generate adversarial examples for models of code based on the semantic equivalent program transformations. Their experimental results demonstrated that QMDP generates adversarial examples and enhances the robustness of the source code classification models by over 44%.
@@
Pour et al. [114] developed a search-based adversarial test generation framework to measure the robustness of the neural source code embedding methods (i.e., Code2vec, Code2seq, and CodeBERT). Their experimental results show that the generated adversarial examples can, on average, decrease the performance of these embedding methods from 5.41% to 9.58%. Applis et al. [115] proposed a testing framework: LAMPION, to assess the robustness of the ML-based program analysis tools adopting metamorphic program transformations. Zhou et al. [26] studied the robustness of the code comment generation tasks in adversarial settings. They proposed ACCENT to generate adversarial examples by substituting identifiers in code snippets to generate syntactically correct and semantically similar ones to original code snippets. Recently, several works have focused on assessing the robustness of pre-trained code models under adversarial attacks [25, 31, 30, 24]. Yang et al. [25] performed a natural attack on the pre-trained code models. They transformed the code snippet, preserving the original inputs’ operational and natural semantics. Bielik et al. [22], and Springer et al. [23] both evaluated the robustness of models of code in adversarial settings. Schuster et al. [116] demonstrated that carefully designed adversarial files attached to the training corpus make the code completion model vulnerable.